Grouper Corporation manufactures a single product. Monthly production costs incurred in the manufacturing process are shown below for the production of 3,800 units. The utilities and maintenance costs are mixed costs. The fixed portions of these costs are $300 and $200, respectively.

Production in Units

3,800

Production Costs

Direct materials $9,500
Direct labor 30,400
Utilities 1,820
Property taxes 1,500
Indirect labor 5,700
Supervisory salaries 1,900
Maintenance 1,340
Depreciation 3,500
Identify the above costs as variable, fixed, or mixed.

Cost

Direct materials

FixedMixedVariable

Direct labor

FixedMixedVariable

Utilities

FixedMixedVariable

Property taxes

FixedMixedVariable

Indirect labor

FixedMixedVariable

Supervisory salaries

FixedMixedVariable

Maintenance

FixedMixedVariable

Depreciation

FixedMixedVariable

LINK TO TEXT

Calculate the expected costs when production is 5,800 units.
Cost to produce 5,800 units $
Answer #1

Answer:-

Identifying the above costs as Fixed/Mixed/Variable:

Costs Fixed/Mixed/Variable
Direct materials Variable
Direct labor Variable
Utilities Mixed
Property taxes Fixed  
Indirect labor Variable
Supervisory salaries Fixed
Maintenance Mixed
Depreciation Fixed

Calculation of expected costs when production is 5800 units:

(Amount in $)
Cost for 3800 units Cost for 5800 units
A B C=A/3800 D=C*5800 E F
Costs Variable Fixed Variable cost per unit Variable Fixed Total cost
Direct materials 9500 2.50 14500 14500
Direct labor 30400 8.00 46400 46400
Utilities 1520 300 0.40 2320 300 2620
Property taxes 1500 0.00 0 1500 1500
Indirect labor 5700 1.50 8700 8700
Supervisory salaries 1900 0.00 0 1900 1900
Maintenance 1140 200 0.30 1740 200 1940
Depreciation 3500 0.00 0 3500 3500
Total Cost 81060

Note:Property taxes,supervisory salaries,Depreciation part of utilities($300) and maintenance($200) costs need to considered as fixed and they will not change with the change in production units.

Direct materials,Direct labor,indirect labor,Part of utilities($1520) and maintenance($1140) costs need to be considered as variable and they will change with the change in production units.